# Fisher information flow in artificial neural networks

### v1

@@ Appendix G Testing our methods on data following a correlated Lognormal distribution
We train here ANNs on a data set following a correlated Lognormal distribution. Lognormal distributions apply to scenarios very different from our imaging experiment, such as earth and environmental science [81]. The Lognormal distribution is skewed and has a right tail that decays slowly thus allowing for “outliers”, resulting in a harder parameter estimation task.
@@
Our procedure for generating the data is the following: we sample Gaussian data analogously to Appendix C, with a dimension of $d=10$. We now introduce correlations between the components by multiplying each data point with a Gaussian random matrix (with independent and identically distributed entries of 0 mean and a standard deviation of $0.1$). Afterwards, we exponentiate each component of the data. The marginal distribution of each component then follows a Lognormal distribution with the components being dependent of each other. While the data follows now a complicated distribution, we still have access to the exact FI, which is equal to that of the underlying Gaussian distribution, since the transformation from the Gaussian random variable to the Lognormal random variable is bijective and thus preserves the FI. We train an ANN to predict the location parameter of the underlying Gaussian (see Appendix C) and show the flow of FI in Fig. A6.
@@
Figure A6: Fisher information calculated for a generic parameter ($\theta=0$) for data sampled from a correlated Lognormal distribution. The curves show the FI as a function of layer depth, normalized by the FI of the input data set. With increased training of the network (see the color bar for the number of training epochs), the FI in all layers approaches the FI of the input data. The ANN’s layers from input to output consist of [10, 150, 100, 50, 25, 1] nodes.
@@
We observe that after a few epochs of training, the ANN manages to conserve the FI to a good approximation. In Fig. A7 we show the learning curves, i.e., the MSE estimated using our training and validation data set, for an ANN trained on the Lognormal data set.
@@
Figure A7: Normalized MSE as a function of the number of epochs for the data following a correlated Lognormal distribution. The dark blue line denotes the validation loss and the light blue line the training loss. The loss function is chosen as the MSE normalized by multiplication with the FI of the input data. The dashed horizontal line indicates where the loss is equal to 1. The epoch number where the training loss crosses this horizontal line is marked by a vertical dashed line. This line coincides with the epoch number where the validation loss is near its minimum. The ANN’s layers from input to output consist of [10, 1500, 1000, 350, 150, 1] nodes.
@@
The model is much larger compared to that we used to calculate the flow of FI, in order to detect overfitting more easily. We observe that our early stopping criterion based on the training loss allows for a very good estimate of the epoch after which the model starts overfitting. The parameters we used for the neural networks and to calculate the FI are listed in the following. For observing the FI flow and also for our early stopping criterion we generate training data consisting of 20 00020\ 000 data points at $21$ equally spaced parameters in the range of $\theta\in\{-7,7\}$. For calculating the FI we use 150 000150\,000 data points at $3$ positions with a spacing of $\Delta\theta=0.1$.
@@
Table AIII: Parameters used for calculating the FI with the LFI maximisation algorithm for the Lognormal data. Data is extracted at each layer of the ANNs that are trained to observe the FI flow.
@@
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Table AIV: Neural network architectures and training parameters for the ANN employed to observe the flow of FI and early stopping based on FI for the Lognormal data

### v2

@@ Appendix G TESTING OUR METHODS ON DATA FOLLOWING A CORRELATED LOG-NORMAL DISTRIBUTION
We train here ANNs on a dataset following a correlated log-normal distribution. Log-normal distributions apply to scenarios very different from our imaging experiment, such as earth and environmental science [83]. The log-normal distribution is skewed and has a right tail that decays slowly, thus allowing for “outliers”, resulting in a harder parameter estimation task.
@@
Our procedure for generating the data is the following: We sample Gaussian data analogously to Appendix C, with a dimension of $d=10$. We now introduce correlations between the components by multiplying each data point with a Gaussian random matrix (with independent and identically distributed entries of 0 mean and a standard deviation of $0.1$). Afterward, we exponentiate each component of the data. The marginal distribution of each component then follows a log-normal distribution with the components being dependent of each other. While the data follow now a complicated distribution, we still have access to the exact FI, which is equal to that of the underlying Gaussian distribution, since the transformation from the Gaussian random variable to the log-normal random variable is bijective and, thus, preserves the FI. We train an ANN to predict the location parameter of the underlying Gaussian (see Appendix C) and show the flow of FI in Fig. A6.
@@
Figure A6: Fisher information calculated for a generic parameter ($\theta=0$) for data sampled from a correlated log-normal distribution. The curves show the FI as a function of layer depth, normalized by the FI of the input dataset. With increased training of the network (see the color bar for the number of training epochs), the FI in all layers approaches the FI of the input data. The ANN’s layers from input to output consist of [10, 150, 100, 50, 25, 1] nodes.
@@
We observe that, after a few epochs of training, the ANN manages to conserve the FI to a good approximation. In Fig. A7, we show the learning curves, i.e., the MSE estimated using our training and validation dataset, for an ANN trained on the log-normal dataset.
@@
Figure A7: Normalized MSE as a function of the number of epochs for the data following a correlated log-normal distribution. The dark blue line denotes the validation loss and the light blue line the training loss. The loss function is chosen as the MSE normalized by multiplication with the FI of the input data. The dashed horizontal line indicates where the loss is equal to 1. The epoch number where the training loss crosses this horizontal line is marked by a vertical dashed line. This line coincides with the epoch number where the validation loss is near its minimum. The ANN’s layers from input to output consist of [10, 1500, 1000, 350, 150, 1] nodes.
@@
The model is much larger compared to that we used to calculate the flow of FI, in order to detect overfitting more easily. We observe that our early stopping criterion based on the training loss allows for a very good estimate of the epoch after which the model starts overfitting. The parameters we used for the neural networks and to calculate the FI are listed in Tables AIII and AIV. For observing the FI flow and also for our early stopping criterion we generate training data consisting of 20 00020\ 000 data points at $21$ equally spaced parameters in the range of $\theta\in\{-7,7\}$. For calculating the FI we use 150 000150\,000 data points at three positions with a spacing of $\Delta\theta=0.1$.
@@
Table AIII: Parameters used for calculating the FI with the LFI maximization algorithm for the log-normal data. Data are extracted at each layer of the ANNs that are trained to observe the FI flow.
@@
Learning rate $1\times 10^{-4}$ $5\times 10^{-6}$
Table AIV: Neural network architectures and training parameters for the ANN employed to observe the flow of FI and early stopping based on FI for the log-normal data